OpenStudy (anonymous):

Rent increases every year by 4.75%. How many years will it take for the rent to increase by 50%? the answer is 9 years but i dont know how to get that

OpenStudy (anonymous):

Intrest rates are calculated using the formula Pe\[^{rt}\] e is used for continuously compounded interest. for yearly compounded interest at a rate of 4.75% 1.0475\[^{x}\] = 1.5 take ln of both sides to get x * ln(1.0475) = ln (1.5) x = ln (1.5) / ln (1.0475) x = 8.74 years

OpenStudy (anonymous):

is that clear enough? or do you want me to fix the formatting?

OpenStudy (anonymous):

um yeah can you? i'm a little confusedd